# Gauge-Invariant Longitudinal Modes in the Herwig 7 Electroweak Parton Shower

### v1

@@ 3.1 $q\to q^{\prime}V$ splittings
In particular, Eq. (34) makes transparent that the GI completion introduces a genuinely new (1z)(1-z)-suppressed contribution already in the kinematic massless limit (neglecting mass terms in the kernels), originating from the Goldstone sector and absent in the pure Dawson subtraction, while Eq. (33) shows how this term is continuously connected to the massive, ultracollinear regime through the Ward-identity matching.
Figure 1: Spin-averaged quasi-collinear splitting kernels $P(z)$ for representative $q\to q^{\prime}V$ branchings at fixed transverse scale $\tilde{q}=500~\mathrm{GeV}$, comparing the original Dawson longitudinal prescription (red) with the GI completion (blue). Left: $b\to t\,W^{-}$, for which the Ward-identity (Goldstone) matching terms are Yukawa-enhanced through the heavy-quark mass and induce a visible deformation of the longitudinal contribution over a broad range of $z$. Right: $b\to b\,Z^{0}$, where the Goldstone couplings are parametrically suppressed, and the two prescriptions are practically indistinguishable at the level of the inclusive kernel.
Figure 1 illustrates, at the level of the spin-averaged $q\to q^{\prime}V$ kernel, how the GI completion reorganises the longitudinal sector relative to the Dawson prescription while leaving the transverse contribution unchanged. For $b\to t\,W^{-}$ (left), the GI longitudinal amplitude contains, in addition to the finite Dawson remainder, the Ward-identity-mandated Goldstone-matching term, whose interference with the vector remainder and whose pure-Goldstone contribution are both controlled by Yukawa couplings; the resulting longitudinal piece is therefore sensitive to the heavy-quark masses and can shift the inclusive kernel by an $\mathcal{O}(1)$ amount across intermediate $z$, even at $\tilde{q}=500~\mathrm{GeV}$, where pure mass-power effects are otherwise expected to decouple. By contrast, for $b\to b\,Z^{0}$ (right), the associated Goldstone couplings are suppressed by $m_{b}/M_{W}$ and the GI correction is correspondingly negligible, so that the Dawson and GI kernels coincide to very high accuracy once summed over polarisations. In both channels, the strong rise towards $z\to 1$ reflects the endpoint enhancement intrinsic to the quasi-collinear kinematics used by the shower, and the plots therefore provide a clean, channel-by-channel diagnostic of where the longitudinal completion can be phenomenologically relevant before embedding the kernels into Sudakov evolution.
@@ 3.2 $V\to V^{\prime}V^{\prime\prime}$ splittings
with $g_{012}$ the channel-dependent triple-gauge coupling (e.g. $g_{ZWW}=g\,c_{W}$ in the SM). As for $q\to q^{\prime}V$, the Herwig spin-dependent kernel is obtained by contracting $\widehat{H}$ with the parent spin-density matrix $\rho$ and, if desired, retaining a spin-density matrix for the emitted vectors for subsequent evolution/decay Masouminia:2021kne . In the GI prescription, the longitudinal shower state $0_{*}$ is understood as the Dawson-subtracted longitudinal current supplemented by the Ward-identity (Goldstone) completion. Concretely, for $V_{0}\to V_{1}V_{2}$ we implement this at the level of reduced amplitudes as
$\widehat{H}^{\mathrm{GI}}_{\lambda_{0}\lambda_{1}\lambda_{2}}=\widehat{H}^{\mathrm{Dawson}}_{\lambda_{0}\lambda_{1}\lambda_{2}}+c_{G}\Big[\delta_{\lambda_{0}0_{*}}\,\widehat{H}^{(G_{0})}_{\lambda_{1}\lambda_{2}}+\delta_{\lambda_{1}0_{*}}\,\widehat{H}^{(G_{1})}_{\lambda_{0}\lambda_{2}}+\delta_{\lambda_{2}0_{*}}\,\widehat{H}^{(G_{2})}_{\lambda_{0}\lambda_{1}}\Big$], (39)
@@
and analogously $F_{TLL}$, $F_{LTL}$, $F_{LLT}$ and $F_{LLL}$. In an unpolarised shower $\rho$ is diagonal and $\varphi$-averaging removes interference terms between different helicity states; nevertheless, we keep the polarisation-resolved definitions explicit since they are required by the full spin-correlation machinery. The remaining helicity sectors vanish at $\mathcal{O}(\lambda^{0})$ in the quasi-collinear expansion with the shower longitudinal basis $0^{\ast}$.
Table 2 summarises the non-vanishing reduced amplitudes in the aligned frame in the two longitudinal prescriptions. The Dawson picture uses only the subtracted longitudinal current ($0_{*}$) without explicit Goldstone-matching. The GI completion implements the Ward-identity completion of the longitudinal sector in the broken theory. Concretely, the transverse block ($\lambda=\pm$) is unchanged, while entries involving $0_{*}$ differ by fixed mass- and $z$-dependent terms.
@@
$\lambda_{0}$ $\lambda_{1}$ $\lambda_{2}$ $\widehat{H}^{\mathrm{Dawson}}_{\lambda_{0}\lambda_{1}\lambda_{2}}=\widehat{H}^{\mathrm{GI}}_{\lambda_{0}\lambda_{1}\lambda_{2}}$
@@
$\lambda_{0}$ $\lambda_{1}$ $\lambda_{2}$ $\widehat{H}^{\mathrm{Dawson}}_{\lambda_{0}\lambda_{1}\lambda_{2}}$ $\widehat{H}^{\mathrm{GI}}_{\lambda_{0}\lambda_{1}\lambda_{2}}$
@@
Table 2: Non-vanishing reduced helicity amplitudes for $V_{0}\to V_{1}V_{2}$ in the Dawson and GI prescriptions, using the Herwig 7 phase convention $\varphi$ and the dimensionless ratios $\mu_{i}=m_{i}/\sqrt{t}$ and $\kappa$ defined in Eq. (48). Entries common to both prescriptions are grouped in the upper block; differences are confined to the lower block and involve the longitudinal state $0_{*}$.
@@
$F_{TTT}^{\mathrm{Dawson}}(z,\tilde{q}^{2})$ $=F_{TTT}^{\mathrm{GI}}(z,\tilde{q}^{2})$ (49)
@@
By contrast, the sectors with a single longitudinal leg differ between Dawson and GI. For the configuration with $V_{2}$ longitudinal, one finds
$F_{TTL}^{\mathrm{Dawson}}(z,\tilde{q}^{2})$ $=2\,m_{2}^{2}\,\left(\rho_{--}+\rho_{++}\right)\frac{z}{\tilde{q}^{2}\,(1-z)^{3}}\,$, (50)
@@
$F_{TLT}^{\mathrm{Dawson}}(z,\tilde{q}^{2})$ $=2\,m_{1}^{2}\,\left(\rho_{--}+\rho_{++}\right)\frac{1-z}{\tilde{q}^{2}\,z^{3}}\,$, (52)
@@
$F_{LTT}^{\mathrm{Dawson}}(z,\tilde{q}^{2})$ $=4\,m_{0}^{2}\,\rho_{00}\,\frac{1-z}{\tilde{q}^{2}\,z}\,$, (54)
@@
It is often useful to write the difference between the GI and Dawson prescriptions directly. Summing the polarisation-resolved contributions into the spin-summed splitting function $F_{(\mathrm{X})}(z,\tilde{q}^{2})$, the net shift induced by the GI completion can be expressed compactly as
$F_{\mathrm{GI}}^{V\to V^{\prime}V^{\prime\prime}}(z,\tilde{q}^{2})-F_{\mathrm{Dawson}}^{V\to V^{\prime}V^{\prime\prime}}(z,\tilde{q}^{2})$ $=2\;c_{G}^{2}\;(\rho_{--}+\rho_{++})\;\frac{\,(m_{1}^{2}+m_{2}^{2})}{\tilde{q}^{2}\,(1-z)\,z}\,$
@@
Figure 2: Comparison of the analytic $V_{0}\to V_{1}V_{2}$ splitting kernels in the Dawson and GI prescriptions for three representative channels: $W\to WZ$ (left), $Z\to W^{+}W^{-}$ (middle) and $\gamma\to W^{+}W^{-}$ (right). In each case we show the spin-summed kernel $F^{(\mathrm{X})}(z,\tilde{q}^{2})$ as a function of the lightcone momentum fraction $z$ at fixed evolution scale $\tilde{q}$ (with pole masses for the external vectors).
Figure 2 illustrates that the impact of the GI completion in bosonic splittings is highly structured: the Dawson and GI pictures share an identical $TTT$ contribution, but differ whenever exactly one daughter is in the longitudinal shower state $0_{*}$. The net effect is a channel-independent, symmetry-breaking shift proportional to $(m_{1}^{2}+m_{2}^{2})$, Eq. (56), which is enhanced in asymmetric branchings through the 1/[z(1z)]1/[z(1-z)] factor and is suppressed as $\tilde{q}$ becomes large compared to the electroweak scale. This makes $V\to V^{\prime}V^{\prime\prime}$ branchings a particularly clean arena to isolate longitudinal-systematics: any deviation between the two schemes can be traced back to the Ward-identity completion of the single-$L$ sectors, rather than to the transverse dynamics, which is common to both constructions.
@@ 4 Results and Discussion
For both the single-resummed and full-resummed studies in this section, we use Herwig’s internal leading-order matrix elements, and keep the shower configuration identical between longitudinal prescriptions except for the choice of the longitudinal scheme flag (Dawson vs. GI). Unless stated otherwise, we analyse events with the Herwig 7-Rivet interface, running custom Rivet Bierlich:2019rhm ; Bierlich:2024vqo analyses and writing histograms in YODA Buckley:2023xqh format for subsequent post-processing with the standard YODA tools. In the single-resummed tests, we isolate the shower effects by switching off multi-parton interactions, hadronisation, and decays, and by limiting the evolution to a single final-state emission; electroweak gauge bosons are kept stable. In the full-resummed, LHC-like runs, we enable the complete shower evolution (with the process-dependent settings given in the input files), retaining the same PDF choice, electroweak scheme, and kinematic cuts as in the corresponding single-emission setups. All generator input files and Rivet analysis routines used to produce the figures in this section are provided alongside this paper.
@@ 4.1 Single-Resummed Evolution
A further, and qualitatively distinct, validation step is to compare the single-emission resummed predictions obtained in the Dawson and GI prescriptions. Fixed-order comparisons, as already presented for the Dawson implementation in Ref. Masouminia:2021kne , probe the correctness of the quasi-collinear kernels at $\mathcal{O}(\alpha_{\mathrm{EW}})$, but they do not test the way in which those kernels are embedded into a Markovian shower evolution, i.e. the interplay between the helicity-dependent branching operator, the shower kinematics map, and the Sudakov exponentiation which enforces unitarity. By restricting to a configuration in which at most one final-state electroweak emission is generated, one obtains a controlled observable bridge between fixed-order and the fully iterated shower: the emission probability is still governed by the resummed (Sudakov-suppressed) single-branching functional, yet it remains sufficiently simple that differences can be traced directly to the longitudinal construction rather than to higher-multiplicity histories, competing emissions, hadronisation, or decay modelling. This is particularly relevant here because the Dawson and GI prescriptions coincide identically in the transverse sector and differ only through the longitudinal completion implied by broken-theory Ward identities; a single-resummed test therefore isolates whether these longitudinal differences survive the shower embedding in a physically meaningful way, and whether they appear in the expected channels and kinematic regions (notably those sensitive to symmetry-breaking and Yukawa structures) before turning to fully exclusive multi-emission phenomenology.
@@
Figure 3: Single-resummed final-state radiation (SRS) comparison between the Dawson and GI longitudinal prescriptions for $q\to q^{\prime}V$ branchings in a clean $e^{+}e^{-}\to q\bar{q}$ setup at $\sqrt{s}=10~\mathrm{TeV}$ (no hadronisation and no decays; exactly one electroweak gauge boson in the event record). Left column: charged-current emission ($q\to q^{\prime}W^{+}$); right column: neutral-current emission ($q\to qZ^{0}$). Shown are the boson transverse-momentum spectrum $p_{\perp,V}$ (top) and the lightcone momentum fraction proxy $z$ used in the shower mapping (bottom); ratio panels display GI/Dawson.
@@
Figure 3 demonstrates, in a deliberately controlled single-emission Sudakov setting, how the GI completion propagates beyond the level of analytic kernels into a shower-embedded prediction. In both $q\to q^{\prime}W^{+}$ and $q\to qZ^{0}$, the Dawson and GI results remain close over most of the kinematic range, with a smooth, predominantly shape-level scheme dependence that is most visible in the soft/collinear region of the $p_{\perp,V}$ spectrum and in the associated redistribution in the mapped energy-sharing variable $z$. The charged-current channel exhibits the larger response, as expected for longitudinal-sensitive emissions, while the neutral-current case remains more weakly affected overall. The significance of this test is twofold: first, it provides a direct diagnostic that the GI longitudinal completion is implemented consistently within the Herwig kinematics map and unitarised evolution (since the two prescriptions coincide in the transverse sector), and second, it establishes that any phenomenological differences observed in fully exclusive simulations can be interpreted as originating from longitudinal systematics in the electroweak shower rather than from secondary modelling choices such as multi-emission competition, hadronisation, or unstable-particle decays.
@@
It is worth stressing that an apparent inversion between the size of the analytic splitting kernels and the size of the observed shower-level signatures is not only possible but, in an exclusive single-emission study, entirely expected. At the level of the branching operator, one compares local quasi-collinear kernels $P^{(\mathrm{X})}(t,z)$ (with $\mathrm{X}=\mathrm{Dawson},\mathrm{GI}$), whereas the event-level distributions in Fig. 3 are generated by a unitarised Markov evolution and are therefore weighted by Sudakov factors. Schematically, the probability density for producing exactly one final-state EW emission at evolution scale $t$ can be written as
@@
is itself controlled by the integral of the same kernel. Consequently, a prescription (such as the GI completion) which yields a locally larger longitudinal contribution in parts of $(t,z)$ phase space typically also yields a stronger Sudakov suppression, and it can enhance the probability for additional emissions. Since our single-resummed validation explicitly targets a one-boson topology (either by construction through emission limits or, equivalently, through an analysis-level requirement of exactly one EW vector boson in the final state), the accepted sample is an exclusive one-emission ensemble. In such an ensemble it is therefore possible for $P_{\mathrm{GI}}>P_{\mathrm{Dawson}}$ at the kernel level to coexist with a larger observed Dawson signature in a given projection: the increased GI emission probability partly migrates into vetoed higher-multiplicity configurations and is partly compensated by the enhanced Sudakov suppression in Eq. (60). A simple toy analogue is a Poisson process with mean $\lambda$, for which the exclusive one-emission probability $\lambda e^{-\lambda}$ decreases once $\lambda$ is increased beyond unity, despite the underlying rate increasing; the shower realisation is a differential, kinematics-dependent version of this mechanism. For this reason, Fig. 3 should be interpreted as a test of the unitarised embedding of the longitudinal prescriptions (kernel plus Sudakov plus kinematics map), rather than as a direct point-by-point reflection of the relative size of the analytic splitting functions.
@@
For the bosonic branchings, we perform an analogous single-resummed study in a V+V+jet topology, designed to isolate the longitudinal-systematics of the $V\to V^{\prime}V^{\prime\prime}$ kernels in a setting that is still sufficiently exclusive to be interpreted in terms of a single shower history. In practice, care is required to ensure that the analysis-level reconstruction mirrors the Herwig shower kinematics as closely as possible and does not inadvertently bias the comparison. First, the recoil object defining the lightlike reference direction $n^{\mu}$ should be identified in an infrared-safe way (e.g. as the hardest coloured final-state parton or via parton-level jet clustering), since overly restrictive definitions such as selecting only a gluon recoil can remove a large fraction of events and distort the reconstructed $z$ and $p_{\perp}$ spectra. Second, the final-state topology used to tag the splitting channel must be made exclusive (e.g. exactly one $W$ and one $Z$ for $W\to WZ$, or exactly one $W^{+}$ and one $W^{-}$ for $Z/\gamma\to W^{+}W^{-}$) to avoid contamination from additional electroweak radiation, and, in particular, $Z\to W^{+}W^{-}$ and $\gamma\to W^{+}W^{-}$ cannot be separated from the final state alone and must be distinguished at the level of the generated hard process. Finally, the kinematic variables entering the comparison should be reconstructed.
@@
Figure 4: SRS shower comparison between the Dawson and GI longitudinal prescriptions for boson-line splittings in a pppp sample at $\sqrt{s}=13~{\rm TeV}$. Left column: splitting-tagged $W^{\pm}\to W^{\pm}Z^{0}$; right column: splitting-tagged $W^{\pm}\to W^{\pm}\gamma$. Upper row: transverse-momentum spectrum of the parent $W^{\pm}$, $p_{\perp,W}$; lower row: reconstructed lightcone momentum fraction $z$ of the splitting as used in the Herwig kinematics map.

### v2

@@ 3.1 $q\to q^{\prime}V$ splittings
In particular, Eq. (34) makes transparent that the GI completion introduces a genuinely new (1z)(1-z)-suppressed contribution already in the kinematic massless limit (neglecting mass terms in the kernels), originating from the Goldstone sector and absent in the pure SL, while Eq. (33) shows how this term is continuously connected to the massive, ultracollinear regime through the Ward-identity matching.
Figure 1: Spin-averaged quasi-collinear splitting kernels $P(z)$ for representative $q\to q^{\prime}V$ branchings at fixed transverse scale $\tilde{q}=500~\mathrm{GeV}$, comparing the original SL prescription (red) with the GI completion (blue). Left: $b\to t\,W^{-}$, for which the Ward-identity (Goldstone) matching terms are Yukawa-enhanced through the heavy-quark mass and induce a visible deformation of the longitudinal contribution over a broad range of $z$. Right: $b\to b\,Z^{0}$, where the Goldstone couplings are parametrically suppressed, and the two prescriptions are practically indistinguishable at the level of the inclusive kernel.
Figure 1 illustrates, at the level of the spin-averaged $q\to q^{\prime}V$ kernel, how the GI completion reorganises the longitudinal sector relative to the SL prescription while leaving the transverse contribution unchanged. For $b\to t\,W^{-}$ (left), the GI longitudinal amplitude contains, in addition to the finite subtraction remainder, the Ward-identity-mandated Goldstone-matching term, whose interference with the vector remainder and whose pure-Goldstone contribution are both controlled by Yukawa couplings; the resulting longitudinal piece is therefore sensitive to the heavy-quark masses and can shift the inclusive kernel by an $\mathcal{O}(1)$ amount across intermediate $z$, even at $\tilde{q}=500~\mathrm{GeV}$, where pure mass-power effects are otherwise expected to decouple. By contrast, for $b\to b\,Z^{0}$ (right), the associated Goldstone couplings are suppressed by $m_{b}/M_{W}$ and the GI correction is correspondingly negligible, so that the SL and GI kernels coincide to very high accuracy once summed over polarisations. In both channels, the strong rise towards $z\to 1$ reflects the endpoint enhancement intrinsic to the quasi-collinear kinematics used by the shower, and the plots therefore provide a clean, channel-by-channel diagnostic of where the longitudinal completion can be phenomenologically relevant before embedding the kernels into Sudakov evolution.
@@ 3.2 $V\to V^{\prime}V^{\prime\prime}$ splittings
with $g_{012}$ the channel-dependent triple-gauge coupling (e.g. $g_{ZWW}=g\,c_{W}$ in the SM). As for $q\to q^{\prime}V$, the Herwig spin-dependent kernel is obtained by contracting $\widehat{H}$ with the parent spin-density matrix $\rho$ and, if desired, retaining a spin-density matrix for the emitted vectors for subsequent evolution/decay Masouminia:2021kne . In the GI prescription, the longitudinal shower state $0_{*}$ is understood as the subtracted longitudinal current supplemented by the Ward-identity (Goldstone) completion. Concretely, for $V_{0}\to V_{1}V_{2}$ we implement this at the level of reduced amplitudes as
$\widehat{H}^{\mathrm{GI}}_{\lambda_{0}\lambda_{1}\lambda_{2}}=\widehat{H}^{\mathrm{SL}}_{\lambda_{0}\lambda_{1}\lambda_{2}}+c_{G}\Big[\delta_{\lambda_{0}0_{*}}\,\widehat{H}^{(G_{0})}_{\lambda_{1}\lambda_{2}}+\delta_{\lambda_{1}0_{*}}\,\widehat{H}^{(G_{1})}_{\lambda_{0}\lambda_{2}}+\delta_{\lambda_{2}0_{*}}\,\widehat{H}^{(G_{2})}_{\lambda_{0}\lambda_{1}}\Big$], (39)
@@
and analogously $F_{TLL}$, $F_{LTL}$, $F_{LLT}$ and $F_{LLL}$. In an unpolarised shower $\rho$ is diagonal and $\varphi$-averaging removes interference terms between different helicity states; nevertheless, we keep the polarisation-resolved definitions explicit, since they are required by the full spin-correlation machinery. The remaining helicity sectors vanish at $\mathcal{O}(\lambda^{0})$ in the quasi-collinear expansion with the shower longitudinal basis $0^{\ast}$.
Table 2 summarises the non-vanishing reduced amplitudes in the aligned frame in the two longitudinal prescriptions. The SL picture uses only the subtracted longitudinal current ($0_{*}$) without explicit Goldstone-matching. The GI completion implements the Ward-identity completion of the longitudinal sector in the broken theory. Concretely, the transverse block ($\lambda=\pm$) is unchanged, while entries involving $0_{*}$ differ by fixed mass- and $z$-dependent terms.
@@
$\lambda_{0}$ $\lambda_{1}$ $\lambda_{2}$ $\widehat{H}^{\mathrm{SL}}_{\lambda_{0}\lambda_{1}\lambda_{2}}=\widehat{H}^{\mathrm{GI}}_{\lambda_{0}\lambda_{1}\lambda_{2}}$
@@
$\lambda_{0}$ $\lambda_{1}$ $\lambda_{2}$ $\widehat{H}^{\mathrm{SL}}_{\lambda_{0}\lambda_{1}\lambda_{2}}$ $\widehat{H}^{\mathrm{GI}}_{\lambda_{0}\lambda_{1}\lambda_{2}}$
@@
Table 2: Non-vanishing reduced helicity amplitudes for $V_{0}\to V_{1}V_{2}$ in the SL and GI prescriptions, using the Herwig 7 phase convention $\varphi$ and the dimensionless ratios $\mu_{i}=m_{i}/\sqrt{t}$ and $\kappa$ defined in Eq. (48). Entries common to both prescriptions are grouped in the upper block; differences are confined to the lower block and involve the longitudinal state $0_{*}$.
@@
$F_{TTT}^{\mathrm{SL}}(z,\tilde{q}^{2})$ $=F_{TTT}^{\mathrm{GI}}(z,\tilde{q}^{2})$ (49)
@@
By contrast, the sectors with a single longitudinal leg differ between SL and GI. For the configuration with $V_{2}$ longitudinal, one finds
$F_{TTL}^{\mathrm{SL}}(z,\tilde{q}^{2})$ $=2\,m_{2}^{2}\,\left(\rho_{--}+\rho_{++}\right)\frac{z}{\tilde{q}^{2}\,(1-z)^{3}}\,$, (50)
@@
$F_{TLT}^{\mathrm{SL}}(z,\tilde{q}^{2})$ $=2\,m_{1}^{2}\,\left(\rho_{--}+\rho_{++}\right)\frac{1-z}{\tilde{q}^{2}\,z^{3}}\,$, (52)
@@
$F_{LTT}^{\mathrm{SL}}(z,\tilde{q}^{2})$ $=4\,m_{0}^{2}\,\rho_{00}\,\frac{1-z}{\tilde{q}^{2}\,z}\,$, (54)
@@
It is often useful to write the difference between the GI and SL prescriptions directly. Summing the polarisation-resolved contributions into the spin-summed splitting function $F_{(\mathrm{X})}(z,\tilde{q}^{2})$, the net shift induced by the GI completion can be expressed compactly as
$F_{\mathrm{GI}}^{V\to V^{\prime}V^{\prime\prime}}(z,\tilde{q}^{2})-F_{\mathrm{SL}}^{V\to V^{\prime}V^{\prime\prime}}(z,\tilde{q}^{2})$ $=2\;c_{G}^{2}\;(\rho_{--}+\rho_{++})\;\frac{\,(m_{1}^{2}+m_{2}^{2})}{\tilde{q}^{2}\,(1-z)\,z}\,$
@@
Figure 2: Comparison of the analytic $V_{0}\to V_{1}V_{2}$ splitting kernels in the SL and GI prescriptions for three representative channels: $W\to WZ$ (left), $Z\to W^{+}W^{-}$ (middle) and $\gamma\to W^{+}W^{-}$ (right). In each case we show the spin-summed kernel $F^{(\mathrm{X})}(z,\tilde{q}^{2})$ as a function of the lightcone momentum fraction $z$ at fixed evolution scale $\tilde{q}$ (with pole masses for the external vectors).
Figure 2 illustrates that the impact of the GI completion in bosonic splittings is highly structured: the SL and GI pictures share an identical $TTT$ contribution, but differ whenever exactly one daughter is in the longitudinal shower state $0_{*}$. The net effect is a channel-independent, symmetry-breaking shift proportional to $(m_{1}^{2}+m_{2}^{2})$, Eq. (56), which is enhanced in asymmetric branchings through the 1/[z(1z)]1/[z(1-z)] factor and is suppressed as $\tilde{q}$ becomes large compared to the electroweak scale. This makes $V\to V^{\prime}V^{\prime\prime}$ branchings a particularly clean arena to isolate longitudinal systematics: any deviation between the two schemes can be traced back to the Ward-identity completion of the single-$L$ sectors, rather than to the transverse dynamics, which is common to both constructions.
@@ 4 Results and Discussion
For both the single-resummed and full-resummed studies in this section, we use Herwig’s internal leading-order matrix elements, and keep the shower configuration identical between longitudinal prescriptions except for the choice of the longitudinal scheme flag (SL vs. GI). Unless stated otherwise, we analyse events with the Herwig 7-Rivet interface, running custom Rivet Bierlich:2019rhm ; Bierlich:2024vqo analyses and writing histograms in YODA Buckley:2023xqh format for subsequent post-processing with the standard YODA tools. In the single-resummed tests, we isolate the shower effects by switching off multi-parton interactions, hadronisation, and decays, and by limiting the evolution to a single final-state emission; electroweak gauge bosons are kept stable. In the full-resummed, LHC-like runs, we enable the complete shower evolution (with the process-dependent settings given in the input files), retaining the same PDF choice, electroweak scheme, and kinematic cuts as in the corresponding single-emission setups. All generator input files and Rivet analysis routines used to produce the figures in this section are provided alongside this paper.
@@ 4.1 Single-Resummed Evolution
A further, and qualitatively distinct, validation step is to compare the single-emission resummed predictions obtained in the SL and GI prescriptions. Fixed-order comparisons, as already presented for the SL implementation in Ref. Masouminia:2021kne , probe the correctness of the quasi-collinear kernels at $\mathcal{O}(\alpha_{\mathrm{EW}})$, but they do not test the way in which those kernels are embedded into a Markovian shower evolution, i.e. the interplay between the helicity-dependent branching operator, the shower kinematics map, and the Sudakov exponentiation which enforces unitarity. By restricting to a configuration in which at most one final-state electroweak emission is generated, one obtains a controlled observable bridge between fixed-order and the fully iterated shower: the emission probability is still governed by the resummed (Sudakov-suppressed) single-branching functional, yet it remains sufficiently simple that differences can be traced directly to the longitudinal construction rather than to higher-multiplicity histories, competing emissions, hadronisation, or decay modelling. This is particularly relevant here because the SL and GI prescriptions coincide identically in the transverse sector and differ only through the longitudinal completion implied by broken-theory Ward identities; a single-resummed test therefore isolates whether these longitudinal differences survive the shower embedding in a physically meaningful way, and whether they appear in the expected channels and kinematic regions (notably those sensitive to symmetry-breaking and Yukawa structures) before turning to fully exclusive multi-emission phenomenology.
@@
Figure 3: Single-resummed final-state radiation (SRS) comparison between the SL and GI longitudinal prescriptions for $q\to q^{\prime}V$ branchings in a clean $e^{+}e^{-}\to q\bar{q}$ setup at $\sqrt{s}=10~\mathrm{TeV}$ (no hadronisation and no decays; exactly one electroweak gauge boson in the event record). Left column: charged-current emission ($q\to q^{\prime}W^{+}$); right column: neutral-current emission ($q\to qZ^{0}$). Shown are the boson transverse-momentum spectrum $p_{\perp,V}$ (top) and the lightcone momentum fraction proxy $z$ used in the shower mapping (bottom); ratio panels display GI/SL.
@@
Figure 3 demonstrates, in a deliberately controlled single-emission Sudakov setting, how the GI completion propagates beyond the level of analytic kernels into a shower-embedded prediction. In both $q\to q^{\prime}W^{+}$ and $q\to qZ^{0}$, the SL and GI results remain close over most of the kinematic range, with a smooth, predominantly shape-level scheme dependence that is most visible in the soft/collinear region of the $p_{\perp,V}$ spectrum and in the associated redistribution in the mapped energy-sharing variable $z$. The charged-current channel exhibits the larger response, as expected for longitudinal-sensitive emissions, while the neutral-current case remains more weakly affected overall. The significance of this test is twofold: first, it provides a direct diagnostic that the GI longitudinal completion is implemented consistently within the Herwig kinematics map and unitarised evolution (since the two prescriptions coincide in the transverse sector), and second, it establishes that any phenomenological differences observed in fully exclusive simulations can be interpreted as originating from longitudinal systematics in the electroweak shower rather than from secondary modelling choices such as multi-emission competition, hadronisation, or unstable-particle decays.
@@
It is worth stressing that an apparent inversion between the size of the analytic splitting kernels and the size of the observed shower-level signatures is not only possible but, in an exclusive single-emission study, entirely expected. At the level of the branching operator, one compares local quasi-collinear kernels $P^{(\mathrm{X})}(t,z)$ (with $\mathrm{X}=\mathrm{SL},\mathrm{GI}$), whereas the event-level distributions in Fig. 3 are generated by a unitarised Markov evolution and are therefore weighted by Sudakov factors. Schematically, the probability density for producing exactly one final-state EW emission at evolution scale $t$ can be written as
@@
is itself controlled by the integral of the same kernel. Consequently, a prescription (such as the GI completion) which yields a locally larger longitudinal contribution in parts of $(t,z)$ phase space typically also yields a stronger Sudakov suppression, and it can enhance the probability for additional emissions. Since our single-resummed validation explicitly targets a one-boson topology (either by construction through emission limits or, equivalently, through an analysis-level requirement of exactly one EW vector boson in the final state), the accepted sample is an exclusive one-emission ensemble. In such an ensemble it is therefore possible for $P_{\mathrm{GI}}>P_{\mathrm{SL}}$ at the kernel level to coexist with a larger observed SL signature in a given projection: the increased GI emission probability partly migrates into vetoed higher-multiplicity configurations and is partly compensated by the enhanced Sudakov suppression in Eq. (60). A simple toy analogue is a Poisson process with mean $\lambda$, for which the exclusive one-emission probability $\lambda e^{-\lambda}$ decreases once $\lambda$ is increased beyond unity, despite the underlying rate increasing; the shower realisation is a differential, kinematics-dependent version of this mechanism. For this reason, Fig. 3 should be interpreted as a test of the unitarised embedding of the longitudinal prescriptions (kernel plus Sudakov plus kinematics map), rather than as a direct point-by-point reflection of the relative size of the analytic splitting functions.
@@
For the bosonic branchings, we perform an analogous single-resummed study in a V+V+jet topology, designed to isolate the longitudinal systematics of the $V\to V^{\prime}V^{\prime\prime}$ kernels in a setting that is still sufficiently exclusive to be interpreted in terms of a single shower history. In practice, care is required to ensure that the analysis-level reconstruction mirrors the Herwig shower kinematics as closely as possible and does not inadvertently bias the comparison. First, the recoil object defining the lightlike reference direction $n^{\mu}$ should be identified in an infrared-safe way (e.g. as the hardest coloured final-state parton or via parton-level jet clustering), since overly restrictive definitions such as selecting only a gluon recoil can remove a large fraction of events and distort the reconstructed $z$ and $p_{\perp}$ spectra. Second, the final-state topology used to tag the splitting channel must be made exclusive (e.g. exactly one $W$ and one $Z$ for $W\to WZ$, or exactly one $W^{+}$ and one $W^{-}$ for $Z/\gamma\to W^{+}W^{-}$) to avoid contamination from additional electroweak radiation, and, in particular, $Z\to W^{+}W^{-}$ and $\gamma\to W^{+}W^{-}$ cannot be separated from the final state alone and must be distinguished at the level of the generated hard process. Finally, the kinematic variables entering the comparison should be reconstructed.
@@
Figure 4: SRS shower comparison between the SL and GI longitudinal prescriptions for boson-line splittings in a pppp sample at $\sqrt{s}=13~{\rm TeV}$. Left column: splitting-tagged $W^{\pm}\to W^{\pm}Z^{0}$; right column: splitting-tagged $W^{\pm}\to W^{\pm}\gamma$. Upper row: transverse-momentum spectrum of the parent $W^{\pm}$, $p_{\perp,W}$; lower row: reconstructed lightcone momentum fraction $z$ of the splitting as used in the Herwig kinematics map.